OpenStudy (anonymous):

need help solving this problem one canned juice drink is 30% orange juice, another is 5% orange juice, how many liters of each should be mixed together in order to get 25L that is 10% orange juice?

OpenStudy (anonymous):

0.30x+0.05x=(0.10)(25)

OpenStudy (amistre64):

amount(%) = useable info

OpenStudy (anonymous):

there u r

OpenStudy (anonymous):

i think i figured it out but want to be sure

OpenStudy (amistre64):

this week is gonna be hecktic for my online personas :) My night time spot is closing early...

OpenStudy (anonymous):

that sucks

OpenStudy (amistre64):

yeah, gonna have to start watching NCIS again lol

OpenStudy (amistre64):

lokis almost a superhero too now , i wish hed hurry it up lol

OpenStudy (anonymous):

x + y = 25 y = 25 -x 0.30x+0.05( 25 -x)=(0.10)(25) now I think you can solve Subtitution: 0.30x+0.05y=(0.10)(25)

OpenStudy (anonymous):

is the answers 5L of the 30% juice and 20L of the 5%

OpenStudy (amistre64):

Nancy Drew gotta it started for us pretty good

OpenStudy (anonymous):

Thank

OpenStudy (amistre64):

we know the total is to be 25; so y = 25-x seems logical right?

OpenStudy (amistre64):

now we substitute amounts for values and percent and destroy this question....with prejudice lol

OpenStudy (amistre64):

even hot the .30 and .05 down pat ;) alot of people do .5

OpenStudy (anonymous):

0.30x+0.05( 25 -x)=(0.10)(25) 0.30x+1.25 -0.05x = 2.5 0.25x = 2.5 -1.25 x = 1.25/0.25 x = 5 I did answer 5

OpenStudy (amistre64):

x = 5; so ys gotta be what 30?

OpenStudy (anonymous):

5+ y = 25 y = 25 -5 y = 20

OpenStudy (anonymous):

i got y=20L and x=5L

OpenStudy (anonymous):

yes, you right

OpenStudy (amistre64):

yeah, but I did georgia math, same answer different result lol

OpenStudy (anonymous):

i did it by the book used the formula for a simular problam

OpenStudy (amistre64):

i had a forman from georgia, his math sense was dumber than a box of rocks, but somehow he always got the right answers ..... I always said it was his georgia math ;)

OpenStudy (anonymous):

thats funny my husband does construction uses measurements all day but cant seem to teach me fractions

OpenStudy (anonymous):

thats funny my husband does construction uses measurements all day but cant seem to teach me fractions